OpenStudy (anonymous):

Jill has been watching the squirrels that live in her neighborhood. The number of squirrels changes each week. n f(n) 1 8 2 12 3 18 4 27 Which function best shows the relationship between n and f(n)?

OpenStudy (anonymous):

f(n) = 12(1.5)n f(n) = 12(1.5)n-1 f(n) = 8(1.5)n f(n) = 8(1.5)n-1

OpenStudy (anonymous):

@ElonaSushchik

OpenStudy (elonasushchik):

Its D for sure

OpenStudy (elonasushchik):

the other ones dont makes sense

OpenStudy (elonasushchik):

@FEARLESS_JOCEY

OpenStudy (anonymous):

That's what I think too. Thanks you are a big help!

OpenStudy (elonasushchik):

you are welcome!! <3

OpenStudy (anonymous):

I have one more question, can you help me with it?

OpenStudy (elonasushchik):

ya sure but in a new tab